OpenStudy (peterg2001):

need help now will medal @kla01 @thebro55 @CritiC

OpenStudy (peterg2001):

One of your classmates was absent on the day that you learned how to write and solve an equation from a contextual situation. Your teacher has asked you to explain the four steps and their application to the problem that follows. Natasha stole three fewer bases than twice the number Elena stole. If Natasha stole 15 bases, how many did Elena steal? Using the four steps from this lesson, write and solve an equation to find the number of bases Elena stole.

OpenStudy (anonymous):

dude i got this from yahoo. see if it helps. Since I don't know what the four steps are, I'll just do it my way Now I'm also not sure if you need to use 1 or 2 variables. -------------------- Method 1: One variable. Let x = number of bases Elena stole Then number of bases Natasha store = 2x - 3 Natasha stole 15 bases: 2x - 3 = 15 How many did Elena steal? Solve for x: 2x - 3 = 15 2x = 18 x = 9 Elena stole 9 bases. -------------------- Method 2. Two variables Let N = number of bases Natasha stole Let E = number of bases Elena stole N = 2E - 3 2a) Find E when N = 15 15 = 2E - 3 2E = 15 + 3 2E = 18 E = 9 2b) Rewrite equation (by solving for E), then find E when N = 15 N = 2E - 3 2E = N + 3 E = (N + 3) / 2 E = (15+3)/2 = 18/2 = 9 Elena stole 9 bases ------------------- As you can see, there are a number of ways to find solution. Without telling me what the mysterious four steps are, that's the best I can do. Source(s): Mind reading abilities not up to par today.

OpenStudy (kla01):

@thebro55 is right. It would be nice to know what the four steps are though.

OpenStudy (anonymous):

try this link. https://answers.yahoo.com/question/index?qid=20101001113651AAMFeaG

OpenStudy (anonymous):

You do 15 (Natascha's Number) + 3, then u divide by 2 :P Try to make that 4 steps